# iFANnpp: Nuclear Power Plant Digital Twin for Robots and Autonomous Intelligence

### v2

@@ VI-B Memory Tests
The Linux top command was used to monitor the memory and CPU usage for each operation. This command was run continuously for 60 seconds, covering the duration during which all profile runs for a specific operation were executed. The instantaneous memory and CPU usage was recorded every second. Both the mean and maximum values for memory and CPU usage were recorded. The memory and CPU usage metrics for each experiment are summarized in the second and third columns of Table I.
To establish a baseline, the usage of memory and CPU was also measured during the idle operation of the iFANnpp Unreal Engine level. In this scenario, no additional operations were performed, providing a reference for comparison with the resource consumption of other tests.
@@ VI-C Functional Tests
For functional testing, unit tests were conducted for each operation to ensure that all developed features functioned as intended. Each unit test verified the correctness and stability of an individual feature. The task for each feature is described in the previous section and tested accordingly. The unit test for each feature starts by launching the project in Unreal Engine. A task operation was executed ten times consecutively. Subsequently, the project was closed and reopened, and the operation was executed ten more times. This cycle was repeated until the operation was performed a total of 100 times. During the development phase, certain features, such as the top-down implementation and VR functionality, caused the project to crash due to bugs that were subsequently identified and resolved.
As shown in I, all tests were ultimately successfully passed, demonstrating the robustness and reliability of the features implemented for research purposes. The consistency of the results across repeated executions underscores the correctness of the implementations and ensures their readiness for practical application in research settings.
@@ VII Conclusion
iFANnpp, a digital twin for nuclear power plants for robots and autonomous intelligence, models a comprehensive nuclear power plant with real-time bidirectional updates with GPWR, the high-fidelity physics simulator for a nuclear power plant. This setup enables researchers to validate robot planning and control methods at a larger scale. The digital twin models the primary physical components of a pressurized water reactor, such as the turbine hall, cooling towers, water sources, and power output area. A Python-AI bridge is provided to enable robotics research areas in numerous areas, incorporating user-friendly access to a range of AI libraries and easier accessibility for users.
@@
The digital twin offers features for designing and testing robot algorithms within nuclear power plants. It supports multiple variants of mobile robots, including wheeled, bipedal, quadruped, and aerial models. Additionally, the digital twin incorporates VR capabilities with controller support, providing users with an immersive environment for more realistic testing and experimentation. For major areas of robotics research, such as task planning and perception, this digital twin also integrates features like navigation utilities, thermal vision, and real-time top-down perspectives.
@@
Various experimental scenarios were conducted to validate the platform’s functionality and reliability. A swarm of robots was created and manually controlled to explore possibilities for multi-robot research. Teleoperation using VR was also investigated, demonstrating the system’s ability to enhance realism in robot control and interaction with nuclear power plant components. The ability to manipulate physical plant components from the Unreal Engine environment and propagate that information to the underlying GPWR simulator adds another level of authenticity to the platform. A navigation task for a robot was investigated using the AI bridge. This scenario validated iFANnpp as a testbed for artificial intelligence applications. RL was employed for a simple robot navigation task, demonstrating the dual advantage of using both tabular data and the digital twin’s realistic visual inputs as data sources for AI models. Additionally, the digital twin underwent comprehensive profiling tests to assess runtime performance. Speed tests, memory usage tests, and unit tests were performed on various platform features, confirming its reliability and effectiveness for robotics research.
@@
Several approaches can be explored to extend this research further. First, the platform can be adapted to model different industrial environments, such as smart factories and manufacturing plants, enabling the testing of robots across diverse industrial environments. Many of the features of the platform can be applied to nuclear power plant operations and a wide range of industrial tasks. Second, the challenges of nuclear power plants can be addressed using advanced AI algorithms, including RL techniques. The Python bridge enables rapid and user-friendly testing and validation of these methods. Finally, additional features relevant to nuclear power plant environments, such as radiation vision and enhanced locomotion capabilities to navigate diverse terrains, can be developed to expand the functionality of the platform further.
@@
iFANnpp holds significant potential for advancing robotics research in nuclear power plants. Its integration of realistic simulations, diverse features, and AI capabilities opens the way for the future of autonomous systems. The platform’s reliability establish it as a robust testbed, supporting further exploration and development in robotics research within the nuclear industry.

### v3

@@ VI-B Memory Tests
The Linux top command was used to monitor the memory and CPU usage for each operation. This command was run continuously for 60 seconds, covering the duration during which all profile runs for a specific operation were executed. Instantaneous memory and CPU usage were recorded every second. Both the mean and maximum values for memory and CPU usage were recorded. The memory and CPU usage metrics for each experiment are summarized in the second and third columns of Table I.
To establish a baseline, the usage of memory and CPU was also measured during the idle state of iFANnpp. In this scenario, no additional operations were performed, providing a reference for comparison with the resource consumption of other tests.
The baseline test consumed roughly 3.37GB of RAM with %108.75 CPU usage. No operation exceeded 5GB of RAM or the use of 5 CPU cores. Therefore the resources required to perform these core feature operations with iFANnpp are accessible to the average researcher.
@@ VI-C Functional Tests
For functional testing, unit tests were conducted for each operation to ensure that all developed features functioned as intended. Each unit test verified the correctness and stability of an individual feature. The task for each feature is described in the previous section and tested accordingly. The unit test for each feature starts by launching the project in UE5. A task operation was executed ten times consecutively. The project was closed and reopened, and the operation was executed ten more times. This cycle was repeated until the operation reached a total of 100 times.
As shown in Table I, all tests were successfully passed. The result demonstrates the robustness and reliability of the digital twin and its features implemented. The perfect result across repeated executions underscores the correctness of the implementations and ensures their readiness for practical application in research settings.
@@ VII Conclusion
iFANnpp, a digital twin for nuclear power plants for robots and autonomous intelligence, models a comprehensive nuclear power plant with real-time bidirectional updates with GPWR, the high-fidelity physics simulator for a nuclear power plant. This setup enables researchers to validate robot planning and control methods at a larger scale. The digital twin models the major physical components of a PWR, such as the turbine hall, cooling towers, water sources, and power output area. A Python-AI bridge is provided to enable robotics and related research areas, incorporating user-friendly access to a range of AI libraries and easier accessibility for users.
@@
The digital twin offers features for designing and testing robot algorithms within nuclear power plants. It supports multiple mobile robots, including wheeled, bipedal, quadruped, and aerial robots. Additionally, the digital twin incorporates the VR feature with controller support, providing users with an immersive environment. For major areas of robotics research, such as task planning and perception, this digital twin also integrates features like navigation utilities, thermal vision, and real-time top-down perspectives.
@@
Various experimental scenarios were conducted to validate the platform’s functionality and reliability. A swarm of robots was created and manually controlled to explore possibilities for multi-robot research. Teleoperation using VR was also investigated, demonstrating the system’s ability to enhance realism in robot control and interaction with nuclear power plant components. Core robot navigation operations were validated, which can be used for RL/AI applications. Additionally, the digital twin underwent comprehensive profiling tests to assess runtime performance. Speed tests, memory usage tests, and unit tests were performed on various platform features, confirming its reliability and effectiveness for robotics research.
@@
Several approaches can be explored to extend this research further. First, the platform can be modeled with different industrial environments, such as smart factories and manufacturing plants. Many of the features of the platform can be applied not only to the nuclear power plant operations but also to a wide range of industrial tasks. Second, the challenges of nuclear power plants can be addressed using advanced AI algorithms, including RL techniques. The Python bridge enables rapid and user-friendly testing and validation of these methods. Finally, additional features relevant to nuclear power plant environments, such as radiation vision and enhanced locomotion capabilities to navigate diverse terrains, can be developed to expand the functionality of the platform further.
@@
iFANnpp holds significant potential for advancing robotics research in nuclear power plants. Its integration of realistic simulations, diverse features, and AI model training capabilities opens the way for the future of autonomous systems. The platform’s reliability establishes it as a robust testbed, supporting further exploration and development in robotics research within the nuclear industry.